# Exactly solvable Schrödinger operators related to the hypergeometric equation

### v1

@@ 7.7 Hyperboloid ${\mathbb{H}}^{p-1,p}$ in complex coordinates
Appendix A Identities for the hypergeometric function
@@
In this section, we review identities for the hypergeometric and Gegenbauer functions with the emphasis on the connection formulas and Kummer’s table. There are many reviews available including [WW] and [NIST1]. Our presentation is close to [D1], [D2].
A.1 Kummer’s table
@@
Recall that the hypergeometric equation is given by the operator ${\cal F}(a,b;c;z,\partial_{z})$ defined in (2.1), and the hypergeometric function with Olver’s normalization $\mathbf{F}(a,b;c;z)$ is defined in (2.2). In this section, for brevity and transparency, we change the notation following [D1] and [D2], writing
$\mathbf{F}_{\alpha,\beta,\mu}(z):=\mathbf{F}\left(\frac{1+\alpha+\beta+\mu}{2},\frac{1+\alpha+\beta-\mu}{2};1+\alpha;z\right).$ (A.1)
@@
$\mathbf{F}_{\alpha,\beta,\mu}(z)=\mathbf{F}_{\alpha,\beta,-\mu}(z).$ (A.2)
The following 6 functions form a set of standard solutions of the hypergeometric equation. Each of the solutions can be expressed in 4 ways (actually, $4\times 2$ ways if we include the trivial identity (A.2)). This yields $6\times 4=24$ expressions usually called Kummer’s table:
$\mathbf{F}_{\alpha,\beta,\mu}(z)$ $=(1-z)^{\frac{-1-\alpha-\beta+\mu}{2}}\mathbf{F}_{\alpha,-\mu,-\beta}(\frac{z}{z-1})$
$=(1-z)^{\frac{-1-\alpha-\beta-\mu}{2}}\mathbf{F}_{\alpha,\mu,\beta}(\frac{z}{z-1})$ (A.3)
@@
$(-z)^{\alpha}\mathbf{F}_{-\alpha,\beta,-\mu}(z)$ $=(-z)^{\alpha}(1-z)^{\frac{-1+\alpha-\beta-\mu}{2}}\mathbf{F}_{-\alpha,\mu,-\beta}(\frac{z}{z-1})$
$=(-z)^{\alpha}(1-z)^{\frac{-1+\alpha-\beta+\mu}{2}}\mathbf{F}_{-\alpha,-\mu,\beta}(\frac{z}{z-1})$ (A.4)
@@
$(-z)^{\frac{-1-\alpha-\beta+\mu}{2}}\mathbf{F}_{-\mu,\beta,-\alpha}(z^{-1})$ $=(1-z)^{\frac{-1-\alpha-\beta+\mu}{2}}\mathbf{F}_{-\mu,\alpha,-\beta}\left(\frac{1}{1-z}\right)$
$=(-z)^{-\alpha}(1-z)^{\frac{-1+\alpha-\beta-\mu}{2}}\mathbf{F}_{-\mu,-\alpha,\beta}\left(\frac{1}{1-z}\right)$ (A.5)
@@
$(-z)^{\frac{-1-\alpha-\beta-\mu}{2}}\mathbf{F}_{\mu,\beta,\alpha}(z^{-1})$ $=(1-z)^{\frac{-1-\alpha-\beta-\mu}{2}}\mathbf{F}_{\mu,-\alpha,-\beta}\left(\frac{1}{1-z}\right)$
$=(-z)^{-\alpha}(1-z)^{\frac{-1+\alpha-\beta+\mu}{2}}\mathbf{F}_{\mu,\alpha,\beta}\left(\frac{1}{1-z}\right)$ (A.6)
@@
$\mathbf{F}_{\beta,\alpha,\mu}(1-z)$ $=(-z)^{\frac{-1-\alpha-\beta+\mu}{2}}\mathbf{F}_{\beta,-\mu,-\alpha}(1-\frac{1}{z})$
$=(-z)^{\frac{-1-\alpha-\beta-\mu}{2}}\mathbf{F}_{\beta,\mu,\alpha}(1-\frac{1}{z})$ (A.7)
@@
$(1-z)^{-\beta}\mathbf{F}_{-\beta,-\mu,-\alpha}(1-z)$ $=(-z)^{\frac{-1-\alpha-\beta+\mu}{2}}\mathbf{F}_{\beta,-\mu,-\alpha}(1-\frac{1}{z})$
$=(-z)^{\frac{-1-\alpha-\beta-\mu}{2}}\mathbf{F}_{\beta,\mu,\alpha}(1-\frac{1}{z})$ (A.8)
@@ A.2 Connection formulas
$\mathbf{F}_{\beta,\alpha,\mu}(1-z)$ (A.9)
$=\frac{\pi~\mathbf{F}_{\alpha,\beta,\mu}(z)}{\sin(-\pi\alpha)\Gamma\left(\frac{1-\alpha+\beta-\mu}{2}\right)\Gamma\left(\frac{1-\alpha+\beta+\mu}{2}\right)}+\frac{\pi~z^{-\alpha}~\mathbf{F}_{-\alpha,\beta,-\mu}(z)}{\sin(\pi\alpha)\Gamma\left(\frac{1+\alpha+\beta+\mu}{2}\right)\Gamma\left(\frac{1+\alpha+\beta-\mu}{2}\right)},$
$(1-z)^{-\beta}\mathbf{F}_{-\beta,\alpha,-\mu}(1-z)$ (A.10)
$=\frac{\pi~\mathbf{F}_{\alpha,\beta,\mu}(z)}{\sin(-\pi\alpha)\Gamma\left(\frac{1-\alpha-\beta-\mu}{2}\right)\Gamma\left(\frac{1-\alpha-\beta-\mu}{2}\right)}+\frac{\pi~z^{-\alpha}~\mathbf{F}_{-\alpha,\beta,-\mu}(z)}{\sin(\pi\alpha)\Gamma\left(\frac{1+\alpha-\beta-\mu}{2}\right)\Gamma\left(\frac{1+\alpha-\beta+\mu}{2}\right)},$
$\mathbf{F}_{\alpha,\beta,\mu}(z)$ (A.11)
$=\frac{\pi~\mathbf{F}_{\beta,\alpha,\mu}(1-z)}{\sin(-\pi\beta)\Gamma\left(\frac{1+\alpha-\beta-\mu}{2}\right)\Gamma\left(\frac{1+\alpha-\beta+\mu}{2}\right)}+\frac{\pi~(1-z)^{-\beta}~\mathbf{F}_{-\beta,\alpha,-\mu}(1-z)}{\sin(\pi\beta)\Gamma\left(\frac{1+\alpha+\beta+\mu}{2}\right)\Gamma\left(\frac{1+\alpha+\beta-\mu}{2}\right)},$
$z^{-\alpha}\mathbf{F}_{-\alpha,\beta,\mu}(z)$ (A.12)
$=\frac{\pi~\mathbf{F}_{\beta,\alpha,\mu}(z)}{\sin(-\pi\beta)\Gamma\left(\frac{1-\alpha-\beta-\mu}{2}\right)\Gamma\left(\frac{1-\alpha-\beta+\mu}{2}\right)}+\frac{\pi~(1-z)^{-\beta}~\mathbf{F}_{-\beta,\alpha,-\mu}(1-z)}{\sin(\pi\beta)\Gamma\left(\frac{1-\alpha+\beta-\mu}{2}\right)\Gamma\left(\frac{1-\alpha+\beta+\mu}{2}\right)}.$
@@
$(-z)^{\frac{-1-\alpha-\beta-\mu}{2}}\mathbf{F}_{\mu,\beta,\alpha}\left(z^{-1}\right)$ (A.13)
$=\frac{\pi~\mathbf{F}_{\alpha,\beta,\mu}(z)}{\sin(-\pi\alpha)\Gamma\left(\frac{1-\alpha-\beta+\mu}{2}\right)\Gamma\left(\frac{1-\alpha+\beta+\mu}{2}\right)}+\frac{\pi~(-z)^{-\alpha}\mathbf{F}_{-\alpha,\beta,\mu}(z)}{\sin(\pi\alpha)\Gamma\left(\frac{1+\alpha-\beta+\mu}{2}\right)\Gamma\left(\frac{1+\alpha+\beta+\mu}{2}\right)};$
$(-z)^{\frac{-1-\alpha-\beta+\mu}{2}}\mathbf{F}_{-\mu,\beta,\alpha}\left(z^{-1}\right)$ (A.14)
$=\frac{\pi~\mathbf{F}_{\alpha,\beta,\mu}(z)}{\sin(-\pi\alpha)\Gamma\left(\frac{1-\alpha-\beta-\mu}{2}\right)\Gamma\left(\frac{1-\alpha+\beta-\mu}{2}\right)}+\frac{\pi~(-z)^{-\alpha}~\mathbf{F}_{-\alpha,\beta,\mu}(z)}{\sin(\pi\alpha)\Gamma\left(\frac{1+\alpha-\beta-\mu}{2}\right)\Gamma\left(\frac{1+\alpha+\beta-\mu}{2}\right)}.$
$\mathbf{F}_{\alpha,\beta,\mu}(z)$ (A.15)
$=\frac{\pi(-z)^{\frac{-1-\alpha-\beta-\mu}{2}}\mathbf{F}_{\mu,\beta,\alpha}(z^{-1})}{\sin(-\pi\mu)\Gamma\left(\frac{1+\alpha+\beta-\mu}{2}\right)\Gamma\left(\frac{1+\alpha-\beta-\mu}{2}\right)}+\frac{\pi(-z)^{\frac{-1-\alpha-\beta+\mu}{2}}\mathbf{F}_{-\mu,\beta,\alpha}(z^{-1})}{\sin(\pi\mu)\Gamma\left(\frac{1+\alpha+\beta+\mu}{2}\right)\Gamma\left(\frac{1+\alpha-\beta+\mu}{2}\right)},$
$(-z)^{-\alpha}\mathbf{F}_{-\alpha,\beta,\mu}(z)$ (A.16)
$=\frac{\pi(-z)^{\frac{-1-\alpha-\beta-\mu}{2}}\mathbf{F}_{\mu,\beta,\alpha}(z^{-1})}{\sin(-\pi\mu)\Gamma\left(\frac{1-\alpha+\beta-\mu}{2}\right)\Gamma\left(\frac{1-\alpha-\beta-\mu}{2}\right)}+\frac{\pi(-z)^{\frac{-1-\alpha-\beta+\mu}{2}}\mathbf{F}_{-\mu,\beta,\alpha}(z^{-1})}{\sin(\pi\mu)\Gamma\left(\frac{1-\alpha+\beta+\mu}{2}\right)\Gamma\left(\frac{1-\alpha-\beta+\mu}{2}\right)}.$
A.3 Degenerate case

### v2

@@ 7.7 Hyperboloid ${\mathbb{H}}^{p-1,p}$ in complex coordinates
Appendix A Holomorphic families of operators
The concept of a holomorphic family of bounded operators is well-known. Less known is the concept of a holomorphic family of closed operators, which we use throughout our paper. Both are described in the monograph of Kato [Kato], Chap. 7, and also in [DW1]. For the convenience of the reader, we give a concise account of these concepts in this appendix.
In the following, $\Theta$ is an open subset of $\mathbb{C}$}, ${\cal H}_{1},{\cal H}_{2}$ are Hilbert spaces, $B({\cal H}_{1},{\cal H}_{2})$ denotes the space of bounded operators from ${\cal H}_{1}$ to ${\cal H}_{2}$ and $C({\cal H}_{1},{\cal H}_{2})$ the set of closed operators from ${\cal H}_{1}$ to ${\cal H}_{2}$. $\mathcal{D}(H)$ will denote the domain of an operator $H$ and $\rho(H):={\mathbb{C}}\backslash\sigma(H)$ its resolvent set.
There exist several equivalent definitions of a holomorphic family of bounded operators. Here is one of them: Consider a function $\Theta\ni z\mapsto H(z)\in B({\cal H}_{1},{\cal H}_{2})$ We say that this is a holomorphic family of bounded operators if for any $f\in{\cal H}_{2}$, $g\in{\cal H}_{1}$ the function $\Theta\ni z\mapsto(f|H(z)g)\in{\mathbb{C}}$ is holomorphic.
We have the following practical criterion:
Theorem A.1.
Suppose that $\Theta\ni z\mapsto H(z)\in B({\cal H}_{1},{\cal H}_{2})$ is a function uniformly bounded on compact subsets of $\Theta$. Suppose in addition that there exist dense subspaces ${\cal X}_{1}\subset{\cal H}_{1}$, ${\cal X}_{2}\subset{\cal H}_{2}$ such that for any $f\in{\cal X}_{1}$, $g\in{\cal X}_{2}$, the function $\Theta\ni z\mapsto(f|H(z)g)\in{\mathbb{C}}$ is holomorphic. Then $\Theta\ni z\mapsto H(z)$ is a holomorphic family of bounded operators.
There exists also a natural concept of a holomorphic family of closed operators, due to Kato, it is however more subtle and less known. Let us describe its definition.
Consider a function $\Theta\ni z\mapsto H(z)\in C({\cal H}_{1},{\cal H}_{2})$. We say that it is a holomorphic family of closed operators if for each $z_{0}\in\Theta$ there exists a neighborhood $\Theta_{0}$ of $z_{0}$, a Hilbert space $\cal K$} and a holomorphic family of bounded operators $\Theta_{0}\ni z\mapsto A(z)\in B({\cal K},{\cal H}_{1})$ that map bijectively $\cal K$} onto $\mathcal{D}(H(z))$ and
$\Theta_{0}\ni z\mapsto H(z)A(z)\in B({\cal K},{\cal H})$
is a holomorphic family of bounded operators.
We have the following practical criterion, which works for operators having non-empty resolvent sets:
Theorem A.2.
Consider a function $\Theta\ni z\mapsto H(z)\in C({\cal H})$. Suppose in addition that for any $z\in\Theta$ the resolvent set of $H(z)$ is nonempty. If for any $z_{0}\in\Theta$ there exists $\lambda\in{\mathbb{C}}$ and a neighborhood $\Theta_{0}$ of $z_{0}$ such that $\lambda\in\rho(H(z))$ for $z\in\Theta_{0}$ and $z\mapsto(H(z)-\lambda)^{-1}\in B({\cal H})$ is holomorphic on $\Theta_{0}$, then $z\mapsto H(z)$ is a holomorphic family of closed operators.
[…]
Appendix B Riemann equation
B.1 Regular-singular points
Consider an ordinary 2nd order differential equation with meromorphic coefficients:
$\left(b(z)\partial_{z}^{2}+c(z)\partial_{z}+d(z)\right)u(z)=0.$ (B.59)
If $\frac{c(z)}{b(z)}$ or $\frac{d(z)}{b(z)}$ are singular at $z_{0}\in{\mathbb{C}}$, then we say that $z_{0}$ is a singular point of (B.59). Following [WW], we distinguish a special class of singular points:
Definition B.1.
We say that (B.59) has a regular-singular point at $z_{0}$, if $\frac{c(z)}{b(z)}$ has a pole of at most 1st order at $z_{0}$, and $\frac{d(z)}{b(z)}$ has a pole of at most 2nd order at $z_{0}$.
We can rewrite (B.59) as
$\left(p(z)(z-z_{0})^{2}\partial_{z}^{2}+q(z)(z-z_{0})\partial_{z}+r(z)\right)u(z)=0.$ (B.60)
If $p(z_{0})\neq 0$, then $z_{0}$ is regular-singular iff $p,q,r$ are analytic at 0. Define the indicial polynomial of $z_{0}$
$P_{z_{0}}(\lambda):=\lambda(\lambda-1)p(z_{0})+\lambda q(z_{0})+r(z_{0}).$ (B.61)
The roots of $P_{z_{0}}$ are called indices of the regular-singular point $z_{0}$.
The importance of regular-singular points and their indices stems from the Frobenius Method, described in the following theorem [WW, Ince]:
Theorem B.2.
Assume that $p(z_{0})\neq 0$ and $p,q,r$ are holomorphic in an open connected simply connected set $\Omega\subset{\mathbb{C}}$ containing $z_{0}$. Let $\rho\in{\mathbb{C}}$ satisfy
$P_{z_{0}}(\rho)=0,\quad P_{z_{0}}(\rho+n)\neq 0,\quad n=1,2,\dots.$ (B.62)
Then there exists a unique function $\tilde{u}(z)$ holomorphic in $\Omega$, such that $u(z):=(z-z_{0})^{\rho}\tilde{u}(z)$ solves (B.60) and $\tilde{u}(z_{0})=1$.
B.2 Riemann equation
Equation with 3 singular points on the Riemann sphere $\mathbb{C}$$\cup\{\infty\}$, all of them regular-singular, is classified by the positions of its singularities and the values of indices. It is called the Riemann, or sometimes the Papperitz equation, and is described in the following well-known theorem [WW, Ince]. For the Riemann operator we use the notation introduced in [DW], which is inspired by the notation due to Papperitz used in [WW].
Theorem B.3.
  • (1)
    Suppose that we are given a 2nd order differential equation on the Riemann sphere having at most 3 singular points $z_{1},z_{2},z_{3}$, all of them regular-singular, with the following indices
    $z_{1}:\ \rho_{1},{\tilde{\rho}}_{1};\quad z_{2}:\ \rho_{2},{\tilde{\rho}}_{2};\quad z_{3}:\ \rho_{3},{\tilde{\rho}}_{3}.$
    Then the following condition is satisfied:
    $\rho_{1}+{\tilde{\rho}}_{1}+\rho_{2}+{\tilde{\rho}}_{2}+\rho_{3}+{\tilde{\rho}}_{3}=1.$ (B.63)
    Such an equation, normalized to have coefficient $1$ at the 2nd derivative, is for finite $z_{1},z_{2},z_{3}$ given by the operator
    ${\cal P}\left(\begin{array}[]{cccc}z_{1}&z_{2}&z_{3}&\\ \rho_{1}&\rho_{2}&\rho_{3}&z,\partial_{z}\\ {\tilde{\rho}}_{1}&{\tilde{\rho}}_{2}&{\tilde{\rho}}_{3}&\end{array}\right):=\partial_{z}^{2}-\left(\frac{\rho_{1}+{\tilde{\rho}}_{1}-1}{z-z_{1}}+\frac{\rho_{2}+{\tilde{\rho}}_{2}-1}{z-z_{2}}+\frac{\rho_{3}+{\tilde{\rho}}_{3}-1}{z-z_{3}}\right)\partial_{z}$
    $+\frac{\rho_{1}{\tilde{\rho}}_{1}(z_{1}-z_{2})(z_{1}-z_{3})}{(z-z_{1})^{2}(z-z_{2})(z-z_{3})}+\frac{\rho_{2}{\tilde{\rho}}_{2}(z_{2}-z_{3})(z_{2}-z_{1})}{(z-z_{2})^{2}(z-z_{3})(z-z_{1})}+\frac{\rho_{3}{\tilde{\rho}}_{3}(z_{3}-z_{1})(z_{3}-z_{2})}{(z-z_{3})^{2}(z-z_{1})(z-z_{2})},$
    and for $z_{3}=\infty$ by the operator
    ${\cal P}\left(\begin{array}[]{cccc}z_{1}&z_{2}&\infty&\\ \rho_{1}&\rho_{2}&\rho_{3}&z,\partial_{z}\\ {\tilde{\rho}}_{1}&{\tilde{\rho}}_{2}&{\tilde{\rho}}_{3}&\end{array}\right):=\partial_{z}^{2}-\left(\frac{\rho_{1}+{\tilde{\rho}}_{1}-1}{z-z_{1}}+\frac{\rho_{2}+{\tilde{\rho}}_{2}-1}{z-z_{2}}\right)\partial_{z}$ (B.67)
    $+\frac{\rho_{1}{\tilde{\rho}}_{1}(z_{1}-z_{2})}{(z-z_{1})^{2}(z-z_{2})}+\frac{\rho_{2}{\tilde{\rho}}_{2}(z_{2}-z_{1})}{(z-z_{2})^{2}(z-z_{1})}+\frac{\rho_{3}{\tilde{\rho}}_{3}}{(z-z_{1})(z-z_{2})}.$ (B.68)
  • (2)
    With the help of homographies (called also Möbius transformations) we can move around singularities. More precisely, if $z\mapsto w(z)=\frac{az+b}{cz+d}$, where we can assume that adbc=1ad-bc=1, then
    ${\cal P}\left(\begin{array}[]{cccc}w(z_{1})&w(z_{2})&w(z_{3})&\\ \rho_{1}&\rho_{2}&\rho_{3}&w,\partial_{w}\\ {\tilde{\rho}}_{1}&{\tilde{\rho}}_{2}&{\tilde{\rho}}_{3}&\end{array}\right)=(cz+d)^{4}{\cal P}\left(\begin{array}[]{cccc}z_{1}&z_{2}&z_{3}&\\ \rho_{1}&\rho_{2}&\rho_{3}&z,\partial_{z}\\ {\tilde{\rho}}_{1}&{\tilde{\rho}}_{2}&{\tilde{\rho}}_{3}&\end{array}\right).$
  • (3)
    By gauging with powers we can shift the indices:
    $\begin{array}[]{l}(z-z_{1})^{-\lambda}(z-z_{2})^{\lambda}{\cal P}\left(\begin{array}[]{cccc}z_{1}&z_{2}&z_{3}&\\ \rho_{1}&\rho_{2}&\rho_{3}&z,\partial_{z}\\ {\tilde{\rho}}_{1}&{\tilde{\rho}}_{2}&{\tilde{\rho}}_{3}&\end{array}\right)(z-z_{1})^{\lambda}(z-z_{2})^{-\lambda}\\[14.22636pt] ={\cal P}\left(\begin{array}[]{cccc}z_{1}&z_{2}&z_{3}&\\ \rho_{1}-\lambda&\rho_{2}+\lambda&\rho_{3}&z,\partial_{z}\\ {\tilde{\rho}}_{1}-\lambda&{\tilde{\rho}}_{2}+\lambda&{\tilde{\rho}}_{3}&\end{array}\right),\end{array}$ (B.69)
    $\begin{array}[]{l}(z-z_{1})^{-\lambda}{\cal P}\left(\begin{array}[]{cccc}z_{1}&z_{2}&\infty&\\ \rho_{1}&\rho_{2}&\rho_{3}&z,\partial_{z}\\ {\tilde{\rho}}_{1}&{\tilde{\rho}}_{2}&{\tilde{\rho}}_{3}&\end{array}\right)(z-z_{1})^{\lambda}\\[14.22636pt] ={\cal P}\left(\begin{array}[]{cccc}z_{1}&z_{2}&z_{3}&\\ \rho_{1}-\lambda&\rho_{2}&\rho_{3}+\lambda&z,\partial_{z}\\ {\tilde{\rho}}_{1}-\lambda&{\tilde{\rho}}_{2}&{\tilde{\rho}}_{3}+\lambda&\end{array}\right).\end{array}$ (B.70)
B.3 From Riemann equation to hypergeometric equation
The Riemann equation can be simplified. First, by Thm B.3 (2), we can assume that the points $z_{1},z_{2},z_{3}$ are any triplet of distinct points on the Riemann sphere. We choose them to be $0,1,\infty$. Then, by Thm B.3 (3), we can assume that $\rho_{1},\rho_{2}$ are arbitrary numbers. We choose them to be both 0. The sum of remaining indices must be $1$. Hence we can assume that 0 has indices 0, 1c1-c; 11 has indices 0, cabc-a-b; and $\infty$ has indices $a$, $b$. The hypergeometric operator is
${\cal F}(a,b;c;z,\partial_{z})$ $:=$ $z(1-z){\cal P}\left(\begin{array}[]{cccc}0&1&\infty&\\ 0&0&a&z,\partial_{z}\\ 1-c&c-a-b&b&\end{array}\right)$ (B.74)
$=$ $z(1-z)\partial_{z}^{2}+(c-(a+b+1)z)\partial_{z}-ab.$ (B.75)
The hypergeometric function $F(a,b;c;z)$ is the unique solution of
${\cal F}(a,b;c;z,\partial_{z})F(z)=0,\quad F(0)=1.$ (B.76)
Thus it is the solution of the hypergeometric equation given by the Frobenius method (Thm B.2) for the singularity $z=0$ and index 0.
Appendix C Identities for the hypergeometric function
@@
In this section, we review identities for the hypergeometric and Gegenbauer functions with the emphasis on the connection formulas and Kummer’s table. There are many reviews available including [WW, NIST1]. Our presentation is close to [D1, D2].
C.1 Kummer’s table
@@
Recall that the hypergeometric equation is given by the operator ${\cal F}(a,b;c;z,\partial_{z})$ defined in (2.1), and the hypergeometric function with Olver’s normalization $\mathbf{F}(a,b;c;z)$ is defined in (2.2). In this section, for brevity and transparency, we change the notation following [D1, D2], writing
$\mathbf{F}_{\alpha,\beta,\mu}(z):=\mathbf{F}\left(\frac{1+\alpha+\beta+\mu}{2},\frac{1+\alpha+\beta-\mu}{2};1+\alpha;z\right).$ (C.1)
@@
$\mathbf{F}_{\alpha,\beta,\mu}(z)=\mathbf{F}_{\alpha,\beta,-\mu}(z).$ (C.2)
The following 6 functions form a set of standard solutions of the hypergeometric equation. Each of the solutions can be expressed in 4 ways (actually, $4\times 2$ ways if we include the trivial identity (C.2)). This yields $6\times 4=24$ expressions usually called Kummer’s table:
$\mathbf{F}_{\alpha,\beta,\mu}(z)$ $=(1-z)^{\frac{-1-\alpha-\beta+\mu}{2}}\mathbf{F}_{\alpha,-\mu,-\beta}(\frac{z}{z-1})$ (C.3)
$=(1-z)^{\frac{-1-\alpha-\beta-\mu}{2}}\mathbf{F}_{\alpha,\mu,\beta}(\frac{z}{z-1})$
@@
$(-z)^{\alpha}\mathbf{F}_{-\alpha,\beta,-\mu}(z)$ $=(-z)^{\alpha}(1-z)^{\frac{-1+\alpha-\beta-\mu}{2}}\mathbf{F}_{-\alpha,\mu,-\beta}(\frac{z}{z-1})$ (C.4)
$=(-z)^{\alpha}(1-z)^{\frac{-1+\alpha-\beta+\mu}{2}}\mathbf{F}_{-\alpha,-\mu,\beta}(\frac{z}{z-1})$
@@
$(-z)^{\frac{-1-\alpha-\beta+\mu}{2}}\mathbf{F}_{-\mu,\beta,-\alpha}(z^{-1})$ $=(1-z)^{\frac{-1-\alpha-\beta+\mu}{2}}\mathbf{F}_{-\mu,\alpha,-\beta}\left(\frac{1}{1-z}\right)$ (C.5)
$=(-z)^{-\alpha}(1-z)^{\frac{-1+\alpha-\beta-\mu}{2}}\mathbf{F}_{-\mu,-\alpha,\beta}\left(\frac{1}{1-z}\right)$
@@
$(-z)^{\frac{-1-\alpha-\beta-\mu}{2}}\mathbf{F}_{\mu,\beta,\alpha}(z^{-1})$ $=(1-z)^{\frac{-1-\alpha-\beta-\mu}{2}}\mathbf{F}_{\mu,-\alpha,-\beta}\left(\frac{1}{1-z}\right)$ (C.6)
$=(-z)^{-\alpha}(1-z)^{\frac{-1+\alpha-\beta+\mu}{2}}\mathbf{F}_{\mu,\alpha,\beta}\left(\frac{1}{1-z}\right)$
@@
$\mathbf{F}_{\beta,\alpha,\mu}(1-z)$ $=(-z)^{\frac{-1-\alpha-\beta+\mu}{2}}\mathbf{F}_{\beta,-\mu,-\alpha}(1-\frac{1}{z})$ (C.7)
$=(-z)^{\frac{-1-\alpha-\beta-\mu}{2}}\mathbf{F}_{\beta,\mu,\alpha}(1-\frac{1}{z})$
@@
$(1-z)^{-\beta}\mathbf{F}_{-\beta,-\mu,-\alpha}(1-z)$ $=(-z)^{\frac{-1-\alpha-\beta+\mu}{2}}\mathbf{F}_{\beta,-\mu,-\alpha}(1-\frac{1}{z})$ (C.8)
$=(-z)^{\frac{-1-\alpha-\beta-\mu}{2}}\mathbf{F}_{\beta,\mu,\alpha}(1-\frac{1}{z})$
@@ C.2 Connection formulas
$\mathbf{F}_{\beta,\alpha,\mu}(1-z)$
$=\frac{\pi~\mathbf{F}_{\alpha,\beta,\mu}(z)}{\sin(-\pi\alpha)\Gamma\left(\frac{1-\alpha+\beta-\mu}{2}\right)\Gamma\left(\frac{1-\alpha+\beta+\mu}{2}\right)}+\frac{\pi~z^{-\alpha}~\mathbf{F}_{-\alpha,\beta,-\mu}(z)}{\sin(\pi\alpha)\Gamma\left(\frac{1+\alpha+\beta+\mu}{2}\right)\Gamma\left(\frac{1+\alpha+\beta-\mu}{2}\right)},$ (C.9)
$(1-z)^{-\beta}\mathbf{F}_{-\beta,\alpha,-\mu}(1-z)$
$=\frac{\pi~\mathbf{F}_{\alpha,\beta,\mu}(z)}{\sin(-\pi\alpha)\Gamma\left(\frac{1-\alpha-\beta-\mu}{2}\right)\Gamma\left(\frac{1-\alpha-\beta-\mu}{2}\right)}+\frac{\pi~z^{-\alpha}~\mathbf{F}_{-\alpha,\beta,-\mu}(z)}{\sin(\pi\alpha)\Gamma\left(\frac{1+\alpha-\beta-\mu}{2}\right)\Gamma\left(\frac{1+\alpha-\beta+\mu}{2}\right)},$ (C.10)
$\mathbf{F}_{\alpha,\beta,\mu}(z)$
$=\frac{\pi~\mathbf{F}_{\beta,\alpha,\mu}(1-z)}{\sin(-\pi\beta)\Gamma\left(\frac{1+\alpha-\beta-\mu}{2}\right)\Gamma\left(\frac{1+\alpha-\beta+\mu}{2}\right)}+\frac{\pi~(1-z)^{-\beta}~\mathbf{F}_{-\beta,\alpha,-\mu}(1-z)}{\sin(\pi\beta)\Gamma\left(\frac{1+\alpha+\beta+\mu}{2}\right)\Gamma\left(\frac{1+\alpha+\beta-\mu}{2}\right)},$ (C.11)
$z^{-\alpha}\mathbf{F}_{-\alpha,\beta,\mu}(z)$
$=\frac{\pi~\mathbf{F}_{\beta,\alpha,\mu}(z)}{\sin(-\pi\beta)\Gamma\left(\frac{1-\alpha-\beta-\mu}{2}\right)\Gamma\left(\frac{1-\alpha-\beta+\mu}{2}\right)}+\frac{\pi~(1-z)^{-\beta}~\mathbf{F}_{-\beta,\alpha,-\mu}(1-z)}{\sin(\pi\beta)\Gamma\left(\frac{1-\alpha+\beta-\mu}{2}\right)\Gamma\left(\frac{1-\alpha+\beta+\mu}{2}\right)}.$ (C.12)
@@
$(-z)^{\frac{-1-\alpha-\beta-\mu}{2}}\mathbf{F}_{\mu,\beta,\alpha}\left(z^{-1}\right)$
$=\frac{\pi~\mathbf{F}_{\alpha,\beta,\mu}(z)}{\sin(-\pi\alpha)\Gamma\left(\frac{1-\alpha-\beta+\mu}{2}\right)\Gamma\left(\frac{1-\alpha+\beta+\mu}{2}\right)}+\frac{\pi~(-z)^{-\alpha}\mathbf{F}_{-\alpha,\beta,\mu}(z)}{\sin(\pi\alpha)\Gamma\left(\frac{1+\alpha-\beta+\mu}{2}\right)\Gamma\left(\frac{1+\alpha+\beta+\mu}{2}\right)};$ (C.13)
$(-z)^{\frac{-1-\alpha-\beta+\mu}{2}}\mathbf{F}_{-\mu,\beta,\alpha}\left(z^{-1}\right)$
$=\frac{\pi~\mathbf{F}_{\alpha,\beta,\mu}(z)}{\sin(-\pi\alpha)\Gamma\left(\frac{1-\alpha-\beta-\mu}{2}\right)\Gamma\left(\frac{1-\alpha+\beta-\mu}{2}\right)}+\frac{\pi~(-z)^{-\alpha}~\mathbf{F}_{-\alpha,\beta,\mu}(z)}{\sin(\pi\alpha)\Gamma\left(\frac{1+\alpha-\beta-\mu}{2}\right)\Gamma\left(\frac{1+\alpha+\beta-\mu}{2}\right)}.$ (C.14)
$\mathbf{F}_{\alpha,\beta,\mu}(z)$
$=\frac{\pi(-z)^{\frac{-1-\alpha-\beta-\mu}{2}}\mathbf{F}_{\mu,\beta,\alpha}(z^{-1})}{\sin(-\pi\mu)\Gamma\left(\frac{1+\alpha+\beta-\mu}{2}\right)\Gamma\left(\frac{1+\alpha-\beta-\mu}{2}\right)}+\frac{\pi(-z)^{\frac{-1-\alpha-\beta+\mu}{2}}\mathbf{F}_{-\mu,\beta,\alpha}(z^{-1})}{\sin(\pi\mu)\Gamma\left(\frac{1+\alpha+\beta+\mu}{2}\right)\Gamma\left(\frac{1+\alpha-\beta+\mu}{2}\right)},$ (C.15)
$(-z)^{-\alpha}\mathbf{F}_{-\alpha,\beta,\mu}(z)$
$=\frac{\pi(-z)^{\frac{-1-\alpha-\beta-\mu}{2}}\mathbf{F}_{\mu,\beta,\alpha}(z^{-1})}{\sin(-\pi\mu)\Gamma\left(\frac{1-\alpha+\beta-\mu}{2}\right)\Gamma\left(\frac{1-\alpha-\beta-\mu}{2}\right)}+\frac{\pi(-z)^{\frac{-1-\alpha-\beta+\mu}{2}}\mathbf{F}_{-\mu,\beta,\alpha}(z^{-1})}{\sin(\pi\mu)\Gamma\left(\frac{1-\alpha+\beta+\mu}{2}\right)\Gamma\left(\frac{1-\alpha-\beta+\mu}{2}\right)}.$ (C.16)
C.3 Degenerate case